Passing through (-3,-5) parallel to the graph of 3x-2y=7
Answer by mananth(16946)

3 x-2y= 7
Find the slope of this line
make y the subject
-2y=-3x +7
Divide by -2
y = 3/ 2x-7/2
Compare this equation with y=mx+b
slope m = 3/ 2
The slope of a line parallel to the above line will be the same
The slope of the required line will be 3/ 2
m= 3/ 2 ,point ( -3 , -5 )
Find b by plugging the values of m & the point in y=mx+b

-5 = -9/ 2 + b
b= - 1/ 2
m= 3/2
Plug value of the slope and b in y = mx +b
The required equation is y =3/2 x- 1/2